OpenStudy (anonymous):

A tank can be filled in 18 hours by Pipe A or in 24 hours by Pipe B. How much time would be needed for both pipes to fill the tank? A) 10 1/3 Hr B)10 9/7 Hr C)10 6/7 Hr D)10 2/7 Hr

OpenStudy (anonymous):

\[\frac{1}{18}+\frac{1}{24}=\frac{1}{x} \]\[x=\frac{72}{7}=10\frac{2}{7} \]